You’re flying a toy plane on a string, and it’s going around at 20.0 m/sec, 10.0 m from you. If it accelerates to a final velocity of 30.0 m/sec in 80.0 seconds, what is its angular acceleration?



1
Expert's answer
2021-10-24T18:27:03-0400

Let's first find the tangential acceleration of the toy plane:


"a_t=\dfrac{v-v_0}{t}=\dfrac{30\ \dfrac{m}{s}-20\ \dfrac{m}{s}}{80\ s}=0.125\ \dfrac{m}{s^2}."

We can find the angular acceleration of the toy plane from the formula:


"a_t=r\alpha,""\alpha=\dfrac{a_t}{r}=\dfrac{0.125\ \dfrac{m}{s^2}}{10\ m}=0.0125\ \dfrac{rad}{s^2}."